(10) Solve y + 6y + 25 y=te-2t with y(0) = 3 and and y(0) = 8. Describe the long term behavior of the system with some dis
Answer #1

Cliven that y+ 69 +284 = t ezt y (0) = 3, y lo] =8 for Homogeneous paut Y+6914254 so the Aumlary equation is leiven By: m²tiss we have formula peut seperti tem wa) = con ven f(0+2) .eve | د ا . --2 D7604211 V1A= as (0-252 +6 (0-2) + 25 :50+4-4D +6WWW. IM H OWUWWWWWWWWWWWWWWWWWWWWWWWWWWWW WW!! 3x289 f + y coj=.8 so find y(t), iss y (t) = el este cosett ez e 3t sinat & t4 Ca = 8+253 462 = 261 :eq=20365 [2= 651 Attoor This y(t) = 151 1034 t + os sinatzt & + (1 - - 12 ) ) Here N e since as ati s